OpenStudy (anonymous):

The ratio of weight on the moon to weight on earth is 1:6. How many pounds would a 144 pound person weigh on the moon?

OpenStudy (gorv):

let weigth on moon = x \[\frac{ weight- on -moon }{ weight-on-earth }=\frac{ x }{ 144 }\]

OpenStudy (gorv):

\[\frac{ 1 }{ 6 }=\frac{ x }{ 144 }\]

OpenStudy (gorv):

multiply bith side by 144

OpenStudy (gorv):

144/6=x x=24